OpenStudy (anonymous):

In a 45-45-90 triangle, the hypotenuse is 18.56 cm. Find the length of the leg. 13.21 cm 13.12 cm 26.25 cm 26.52 cm .

OpenStudy (anonymous):

18.56cos45=13.12 cm

OpenStudy (zarkon):

For a 45-45-90 triangle the lengths of the legs are the same. Therefore, \[\ell^2+\ell^2=h^2\Rightarrow 2\cdot\ell^2=h^2\Rightarrow\sqrt{2}\ell=h\Rightarrow\ell=\frac{h}{\sqrt{2}}\] \[\ell=\frac{18.56}{\sqrt{2}}\approx13.12\]

OpenStudy (anonymous):

thanks so much :)